# Emergence of Goal-Directed Behaviors via Active Inference with Self-Prior

### v1

@@ 1 Introduction
Infants spontaneously exhibit goal-directed behaviors, such as reaching for a sensory stimulus or actively exploring their surroundings, even in the absence of external rewards essential for survival. These behaviors are highly motivated by internal satisfaction and are referred to as intrinsically motivated behaviors (Czikszentmihalyi,, 1990; Ryan & Deci,, 2000). Such behaviors facilitate unsupervised learning without explicit external rewards or punishments, which is known to offer various advantages in early developmental stages (White,, 1959; Gopnik,, 2009; Zaadnoordijk et al.,, 2022; Kanazawa et al.,, 2023). For example, sensory experiences arising from an infant’s spontaneous movements, such as self-touch, not only promote learning about one’s own body (e.g., acquiring body representations) (Hoffmann et al.,, 2017), but also contribute to the formation of an early sense of self (Rochat,, 1998).
@@
Traditional observational studies and neuroscience methodologies have been employed to investigate the mechanisms underlying intrinsically motivated behavioral development (Di Domenico & Ryan,, 2017). In particular, computational modeling has also been utilized to quantitatively interpret and predict behavioral development (Oudeyer & Kaplan,, 2009; Shultz,, 2013). Experiments using computer simulations allow researchers to freely manipulate and control specific variables to investigate various scenarios, making them advantageous for uncovering latent effects and interactions that may be difficult to observe through conventional behavioral studies alone.
@@
Although computational modeling is a useful approach for studying infant behavioral development, many models proposed in modern robotics and machine learning tend to treat intrinsic motivation primarily as an exploratory mechanism for improving reward acquisition efficiency in environments with sparse external rewards (Aubret et al.,, 2023). Indeed, such approaches can enhance reinforcement learning agents’ ability to improve explicit task performance, such as maximizing video game scores. However, they do not sufficiently address how entirely new behaviors emerge when no explicit external reward objective is set, meaning that no performance criterion or information gain is provided.
@@
To address this limitation, we propose a novel computational model of intrinsic motivation based on the expected free energy (Friston,, 2010; Friston et al.,, 2016) (Figure 1). We introduce an internal density model, termed the self-prior, which enables an agent to learn a statistical representation of incoming multimodal sensory signals. When a mismatch between the learned probabilistic model and actual observations is detected, goal-directed behavior emerges without any explicit reward criterion. Specifically, we implement a simulated agent that perceives a stimulus (such as a sticker) on its arm and spontaneously exhibits reaching behavior toward it. In the latter part of this paper, we discuss how the proposed self-prior mechanism may contribute to developmental science, neuroscience, and machine learning, positioning this study as a computational investigation into the early development of intentional behaviors in infants.
@@ 1.1 Free Energy Principle
The free energy principle posits that an agent can learn to perceive and act by minimizing the "surprise" induced by sensory inputs (Friston,, 2010). In the free energy framework, surprise refers to the prediction error that arises when the agent encounters unexpected sensory observations, and reducing this surprise becomes the agent’s primary objective.
At each time step $t$, the agent receives information from the world through its sensory organs. We denote this sensory input by $o_{t}$, referring to it as the “observation.” Although $o_{t}$ is all the agent can directly observe, there may exist underlying factors in the world that generate these observations. Such factors are called hidden states, denoted by $s_{t}$ at time $t$. For example, given a properly learned model, an agent can infer from a tactile pattern on its arm (observation $o_{t}$) that there is a sticker attached to its arm (hidden state $s_{t}$).
Consider an agent that must satisfy certain biological criteria for survival in a stable manner. To do so, the agent must accurately understand its environment, which amounts to inferring $p(o_{t})$. According to the free energy principle, an agent maintains its own probabilistic model of these hidden states because doing so enables it to approximate $p(o_{t})$ and thus make inferences about the world. In practice, we often maximize the log of $p(o_{t})$ (i.e., $\log p(o_{t})$) for computational convenience, which is referred to as maximizing the model evidence. Minimizing the negative term, $-\log p(o_{t})$—known as surprise—is therefore equivalent to maximizing this model evidence.
In reality, calculating $p(o_{t})$ requires marginalization over hidden states (i.e., $p(o_{t},s_{t})$), which is often intractable. The variational free energy principle tackles this by introducing a variational approximation of Bayesian inference, wherein the agent indirectly reduces surprise by minimizing the variational free energy $\mathcal{F}$, an upper bound on $-\log p(o_{t})$. Formally, $\mathcal{F}$ is defined as follows (see (Parr et al.,, 2022) for details):
@@ 1.2 Active Inference and Expected Free Energy
While the free energy principle describes how an agent updates its model to reduce surprise based on current sensory observations, the agent can also directly change its observations through actions $a_{t}$ to make them better align its internal model. This is the essence of active inference, and more concretely, the approximate posterior $q(s_{t})$ is extended to include actions as $q(s_{t},a_{t})=q(a_{t}\mid s_{t})\,q(s_{t})$.
To obtain optimal actions $a_{t}$ for both present and future steps, a method has been proposed that computes expected free energy by considering the expected free energy at future time points (Friston et al.,, 2016). That is, the agent predicts multiple possible future states and selects the trajectory that is expected to minimize free energy most effectively.
However, since future observations cannot be obtained before the corresponding time actually arrives, the agent cannot directly compute the free energy of future states. Therefore, the expected free energy assumes that the agent possesses a preferred prior $\tilde{p}(o_{t})$ over desired observations, which acts as a setpoint. As a result, $p(o_{t},s_{t})$ is extended to include actions and is biased toward preferred observations, expressed as $\tilde{p}(o_{t},s_{t},a_{t})=p(a_{t}\mid s_{t})\,p(s_{t}\mid o_{t},a_{t})\,\tilde{p}(o_{t}\mid a_{t})$. Assuming $q(s_{t}\mid o_{t},a_{t})\approx p(s_{t}\mid o_{t},a_{t})$ and a uniform prior $p(a_{t})$ over actions, the expected free energy $\mathcal{G}$ is defined as follows (see (Millidge et al.,, 2020; Mazzaglia et al.,, 2021) for detailed approximations and derivation):
@@
where $\mathcal{H}$ denotes the information entropy. That is, to reduce surprise at a future time $t$, the agent must (i) increase the probability of preferred observations, (ii) obtain high information gain from those observations, and (iii) maintain diverse actions. As shown above, expected free energy integrates both pragmatic (extrinsic) and epistemic (intrinsic) value into a single expression, thereby naturally resolving the exploration–exploitation dilemma. The extrinsic motivation term is similar to a drive for achieving goals or obtaining rewards necessary for survival, whereas the intrinsic motivation term corresponds to an internal desire to explore new information or reduce uncertainty. Since both are measured in the same information unit (nats), they are automatically balanced within a unified system.
@@ 2 Self-Prior: Acquired Preference to Induce Goal-Directed Behavior
Under the expected free energy, agents do not require scalar rewards to be explicitly defined by the environment. This is because the agent is assumed to possess a distribution over preferred observations, called the preferred prior, which serves as the setpoint for action planning. However, since the preferred prior is typically given in a fixed form, the removal of external reward signals does not imply that the agent has autonomously generated its goal-directed behavior.
@@
By contrast, in this study, we allow the agent to autonomously determine the setpoint for its action planning by additionally learning an internal density model, which we refer to as the self-prior. Through this, the agent spontaneously generates goal-directed behavior to minimize discrepancies between its internal model and the environment, including its own body.
Technically, we derive this by making a simple modification to the standard expected free energy formulation—separating the self-prior term from the original preferred prior term. This enables the extraction of an intrinsic motivation term that maintains preferences formed through experience history, as distinct from the extrinsic motivation term that maintains specific experiences. According to the classification proposed by (Oudeyer & Kaplan,, 2009), this corresponds to internal homeostatic intrinsic motivation, and in particular, is most closely related to what is termed familiarity-driven intrinsic motivation. We further examine in the Discussion section why intrinsic motivation can be extracted from the extrinsic motivation term in the free energy formulation.
@@
To implement the self-prior in the active inference framework, we redefine each preferred observation $o$ in terms of two components: (i) an extrinsic observation $o^{E}$ that must be maintained to satisfy survival requirements, and (ii) an intrinsic observation $o^{I}$ related to the self-prior. Here, $o^{I}$ refers to bodily states or sensations—such as a tactile pattern on the arm—that can vary without threatening the agent’s essential functions. Consequently, we decompose $\tilde{p}(o)$ into $\tilde{p}(o^{E},o^{I})$, and in this study, we assume independence so that $\tilde{p}(o^{E},o^{I})=\tilde{p}(o^{E})\,\tilde{p}(o^{I})$. Thus, the expected free energy incorporating the self-prior becomes:
@@
$\mathcal{G}$ $\approx-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{extrinsic (homeostatic)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (heterostatic)}}-\mathcal{H^{\prime}}$ (3)
$=-\underbrace{\mathbb{E}_{q(o^{E}_{t},o^{I}_{t})}[\log\tilde{p}(o^{E}_{t},o^{I}_{t})]}_{\text{homeostatic}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{heterostatic}}-\mathcal{H^{\prime}}$
@@
$=-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeo-, fixed)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeo-, familiarity)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (hetero-, novelty)}}-\mathcal{H^{\prime}}$
@@
We designed the system such that the agent can learn the self-prior autonomously, by automatically acquiring a mechanism that estimates familiarity based on the frequency of past observations. Details of the implementation are provided in the following section.
@@ 3 Computational Model of Self-Prior
In this section, we introduce the concrete implementation of our model for computational simulations based on the active inference framework incorporating the self-prior. We present separate models for environments with discrete and continuous variables. In the discrete environment, the model is built using small-sized matrices to allow direct inspection of its behavior. For the continuous environment, which requires a high-dimensional variable space, we use deep neural networks with techniques developed in deep reinforcement learning.
3.1 Discrete Model
The discrete model uses categorical distributions as the generative model, handling discrete random variables:

### v2

@@ 1 Introduction
Infants spontaneously exhibit goal-directed behaviors, such as reaching for a sensory stimulus or actively exploring their surroundings, even in the absence of external rewards essential for survival. These behaviors are highly motivated by internal satisfaction and are referred to as intrinsically motivated behaviors (czikszentmihalyi_flow_1990; ryan_intrinsic_2000). Such behaviors facilitate unsupervised learning without explicit external rewards or punishments, which is known to offer various advantages in early developmental stages (white_motivation_1959; gopnik_philosophical_2009; zaadnoordijk_lessons_2022; kanazawa_open-ended_2023). For example, sensory experiences arising from an infant’s spontaneous movements, such as self-touch, not only promote learning about one’s own body (e.g., acquiring body representations) (hoffmann_development_2017), but also contribute to the formation of an early sense of self (rochat_self-perception_1998).
@@
Traditional observational studies and neuroscience methodologies have been employed to investigate the mechanisms underlying intrinsically motivated behavioral development (di_domenico_emerging_2017). In particular, computational modeling has also been utilized to quantitatively interpret and predict behavioral development (oudeyer_what_2009; shultz_computational_2013). Experiments using computer simulations allow researchers to freely manipulate and control specific variables to investigate various scenarios, making them advantageous for uncovering latent effects and interactions that may be difficult to observe through conventional behavioral studies alone.
@@
Although computational modeling is a useful approach for studying infant behavioral development, many models proposed in modern robotics and machine learning tend to treat intrinsic motivation primarily as an exploratory mechanism for improving reward acquisition efficiency in environments with sparse external rewards (aubret_information-theoretic_2023). Indeed, such approaches can enhance reinforcement learning agents’ ability to improve explicit task performance, such as maximizing video game scores. However, they do not sufficiently address how entirely new behaviors emerge when no explicit external reward objective is set, meaning that no performance criterion or information gain is provided.
@@
To address this limitation, we propose a novel computational model of intrinsic motivation based on the expected free energy (friston_free-energy_2010; friston_active_2016) (Figure 1). We introduce an internal density model, termed the self-prior, which enables an agent to learn a statistical representation of incoming multimodal sensory signals. When a mismatch between the learned probabilistic model and actual observations is detected, goal-directed behavior emerges without any explicit reward criterion.
Our approach combines several key features that, to our knowledge, have not been integrated in previous computational models of developmental behavior: (i) the behavioral setpoint is autonomously formed from the agent’s own experience rather than externally predefined; (ii) specific goal-directed behaviors such as reaching emerge independently, rather than serving merely as auxiliary mechanisms for reward exploration; (iii) behaviors persist even without information gain, as long as there is a mismatch with familiar states; and (iv) the mechanism is theoretically integrable with existing extrinsic and intrinsic motivations within the active inference framework. We implement a simulated agent that perceives a stimulus (such as a sticker) on its arm and spontaneously exhibits reaching behavior toward it through the self-prior mechanism. In the latter part of this paper, we discuss how the proposed mechanism may contribute to developmental science, positioning this study as a computational investigation into the early development of intentional behaviors in infants.
@@ 1.1 Free Energy Principle
The free energy principle posits that an agent can learn to perceive and act by minimizing the "surprise" induced by sensory inputs (friston_free-energy_2010). This surprise refers to the prediction error that arises when the agent encounters unexpected sensory observations, and reducing this surprise becomes the agent’s primary objective.
At each time step $t$, the agent receives information from the world through its sensory organs. We denote this sensory input by $o_{t}$, referring to it as the observation. Although $o_{t}$ is all the agent can directly observe, there may exist underlying factors in the world that generate these observations. Such factors are called hidden states, denoted by $s_{t}$ at time $t$. For example, given a properly learned model, an agent can infer from a tactile pattern on its arm (observation $o_{t}$) that there is a sticker attached to its arm (hidden state $s_{t}$).
Consider an agent that must satisfy certain biological criteria for survival in a stable manner. To do so, the agent must accurately understand its environment, which amounts to inferring $p(o_{t})$. According to the free energy principle, an agent maintains its own probabilistic model of these hidden states because doing so enables it to approximate $p(o_{t})$ and thus make inferences about the world. In practice, we often maximize the log of $p(o_{t})$ (i.e., $\log p(o_{t})$) for computational convenience, which is referred to as maximizing the model evidence. Minimizing the negative term, $\log p(o_{t})$, known as surprise, is therefore equivalent to maximizing this model evidence.
In reality, calculating $p(o_{t})$ requires marginalization over hidden states (i.e., $p(o_{t},s_{t})$), which is often intractable. The variational free energy principle tackles this by introducing a variational approximation of Bayesian inference, wherein the agent indirectly reduces surprise by minimizing the variational free energy $\mathcal{F}$, an upper bound on $\log p(o_{t})$. Formally, $\mathcal{F}$ is defined as follows (see (parr_active_2022) for details):
@@ 1.2 Active Inference and Expected Free Energy
While the free energy principle describes how an agent updates its model to reduce surprise based on current sensory observations, the agent can also directly change its observations through actions $a_{t}$ to make them align better with its internal model. This is the essence of active inference, and more concretely, the approximate posterior $q(s_{t})$ is extended to include actions as $q(s_{t},a_{t})=q(a_{t}\mid s_{t})\,q(s_{t})$.
To obtain optimal actions $a_{t}$ for both present and future steps, a method has been proposed that computes expected free energy by considering the expected free energy at future time points (friston_active_2016). That is, the agent predicts multiple possible future states and selects the trajectory that is expected to minimize free energy most effectively.
However, since future observations cannot be obtained before the corresponding time actually arrives, the agent cannot directly compute the free energy of future states. Therefore, the expected free energy assumes that the agent possesses a preferred prior $\tilde{p}(o_{t})$ over desired observations, which acts as a setpoint. As a result, $p(o_{t},s_{t})$ is extended to include actions and is biased toward preferred observations, expressed as $\tilde{p}(o_{t},s_{t},a_{t})=p(a_{t})\,p(s_{t}\mid o_{t})\,\tilde{p}(o_{t})$. Thus, the expected free energy $\mathcal{G}$ is defined as follows (see (millidge_relationship_2020; mazzaglia_contrastive_2021) for details):
@@
where $\mathcal{H}$ denotes the information entropy. That is, to reduce surprise at a future time $t$, the agent must (i) increase the probability of preferred observations, (ii) obtain high information gain from those observations, and (iii) maintain diverse actions. As shown above, expected free energy integrates both pragmatic (extrinsic) and epistemic (intrinsic) value into a single expression, thereby naturally resolving the exploration–exploitation dilemma.
@@ 2 Self-Prior: Acquired Preference to Induce Goal-Directed Behavior
Under the expected free energy, agents do not require scalar rewards to be explicitly defined by the environment. This is because the agent is assumed to possess information about preferred observations as a preferred prior distribution, which serves as the setpoint for action planning.
@@
However, since the preferred prior is typically given in a fixed form, the removal of external reward signals does not imply that the agent has autonomously generated its goal-directed behavior. Therefore, in this study, we allow the agent to autonomously determine the setpoint for its action planning by additionally learning a density model, which we refer to as the self-prior.
The self-prior represents "the probability density over the frequency of observed sensations that the agent has experienced," and is defined as a variable preferred prior that induces behavioral tendencies for the agent to maintain or re-experience familiar states. Unlike fixed preferred priors, this means that models with the same structure can develop different priors depending on different experiences, leading to the emergence of different behaviors. To provide an intuitive understanding of how the self-prior operates, this paper presents the example of spontaneous reaching toward and removal of a sticker attached to one’s own body. This is similar to the experimental setup of (bigelow_development_1986), where a silent toy was placed on the body of a blind infant to examine reaching behavior. Here, we interpret the reason why the infant agent removes a sticker attached to its body as being due to a mismatch between the prior of "oneself without a sticker" and "the currently observed self." Moreover, this prior has been autonomously and gradually developed by the agent through everyday experiences (in which no sticker was attached). Since this self-prior does not contain specific instructions to remove the sticker, it is distinguished from approaches that directly command reaching by providing fixed setpoints.
@@
The implementation of the self-prior in the active inference framework is derived from a simple modification to the standard expected free energy formulation, separating the self-prior term from the original preferred prior term. Specifically, this is achieved by redefining each preferred observation $o$ in terms of two components: (i) an extrinsic observation $o^{E}$ that must be maintained to satisfy survival requirements, and (ii) an intrinsic observation $o^{I}$ related to the self-prior. Here, $o^{I}$ refers to bodily states or sensations such as a tactile pattern on the arm, that can vary without threatening the agent’s essential functions. Consequently, we decompose the preferred prior $\tilde{p}(o)$ into $\tilde{p}(o^{E},o^{I})$, and in this study, we assume independence so that $\tilde{p}(o^{E},o^{I})\approx\tilde{p}(o^{E})\,\tilde{p}(o^{I})$:
Preferred prior: $\tilde{p}(o)=\tilde{p}(o^{E},o^{I})\approx\tilde{p}(o^{E})\,\underbrace{\tilde{p}(o^{I})}_{\mathclap{\textbf{Self-Prior}}}$~~~ (3)
In our approach, we choose to define the self-prior over observations $o$ rather than internal states $s$. This is because, as mentioned earlier, the self-prior can be derived through simple manipulation of the standard expected free energy formulation, enabling natural integration of fixed extrinsic preferences and experience based intrinsic preferences within the same framework. The fact that the self-prior is related to intrinsic motivation yet can emerge from external sensory data may seem counterintuitive. This confusion stems from inconsistent use of "external-extrinsic" and "internal-intrinsic" terminology. To resolve this confusion and rigorously classify intrinsic motivation, we adopt the computational classification of intrinsic motivation proposed by (oudeyer_what_2009):
  • External vs. Internal: Referring to whether the reward computation occurs in the environment or within the agent.
  • Extrinsic vs. Intrinsic: Referring to whether the reward criterion is determined externally or internally.
According to these definitions, any reward provided from the external environment, such as a game score, qualifies as extrinsic motivation. Meanwhile, internal motivation can be either extrinsic or intrinsic. An internal motivation is considered intrinsic if the agent sets its reward criterion according to its internal model (e.g., measuring information gain based on a learned world model). By contrast, an internal yet extrinsic motivation is one in which the agent checks its internal state and computes a reward based on an externally fixed criterion (e.g., "maintaining a battery level of 50%").
Therefore, the key to distinguishing extrinsic and intrinsic motivation is not the source of information, but rather who sets the reward criterion. In this sense, expected free energy is an internal framework since the agent itself computes both extrinsic and intrinsic components. Furthermore, the self-prior proposed in this paper qualifies as internal intrinsic motivation because the agent’s internal model directly calculates and updates the reward criterion.
Additionally, (oudeyer_what_2009) proposed that motivation can also be distinguished as homeostatic vs. heterostatic. For example, if there is an intrinsic motivation to pursue new information acquisition, this is heterostatic, whereas if there is an intrinsic motivation to maintain a familiar state, this is defined as homeostatic.
In active inference, the preferred prior has traditionally been interpreted as a fixed setpoint representing observations essential for survival (e.g., blood sugar level, battery level). These setpoints have been considered extrinsic values since they are externally specified, but we regard this term as strictly representing a homeostatic value characteristic. Therefore, the decomposition in Equation (3) separates the homeostatic value term in Equation (2) into extrinsic and intrinsic terms:
@@
$-\underbrace{\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]}_{\text{homeostatic value}}=-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeostatic)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeostatic)}}$ (4)
Note that unlike Equation (2), where $\mathbb{E}_{q(o_{t})}[\log\tilde{p}(o_{t})]$ is labeled as extrinsic value, Equation (4) labels it as homeostatic. Thus, the expected free energy incorporating the self-prior becomes:
@@
$\mathcal{G}$ $\approx-\underbrace{\mathbb{E}_{q(o^{E}_{t})}[\log\tilde{p}(o^{E}_{t})]}_{\text{extrinsic (homeo-, fixed)}}-\underbrace{\mathbb{E}_{q(o^{I}_{t})}[\log\tilde{p}(o^{I}_{t})]}_{\text{intrinsic (homeo-, familiarity)}}-\underbrace{\mathbb{E}_{q(o_{t})}[D_{\mathrm{KL}}[q(s_{t}|o_{t})\|q(s_{t})]]}_{\text{intrinsic (hetero-, novelty)}}$-$\mathcal{H^{\prime}}$ (5)
@@
In summary, the self-prior in this study aims to "maintain or restore a reference pattern" learned from the agent’s sensory experiences, thus qualifying as internal homeostatic intrinsic motivation. This most closely resembles distributional familiarity motivation (DFM) among the classifications of intrinsic motivation proposed by (oudeyer_what_2009). Table 1 summarizes the positioning of the self-prior within this motivational taxonomy.
Extrinsic Intrinsic
Homeostatic Heterostatic
External Score in video game (N/A) (N/A)
Internal
Gap of sensory data
from fixed value
Gap of sensory data
from self-prior
Information gain
by sensory data
As an alternative approach to implementing the self-prior concept, there also exist approaches that construct preferences over internal states. This is the approach proposed by (sajid_exploration_2021), but it could not explicitly integrate extrinsic and intrinsic motivations. Another approach could consider hierarchical models wherein internal representations learned at higher levels are transformed into self-priors for lower levels. We revisit these potential future extensions in the Discussion section.
@@ 3 Computational Model of Self-Prior
In this section, we introduce the concrete implementation of our model for computational simulations based on the active inference framework incorporating the self-prior. The overall framework is constructed using small-sized matrices in the discrete environment to allow direct inspection of model behavior, and deep neural networks using techniques developed in deep reinforcement learning in the continuous environment.
We implement this as a general-purpose structure applicable to a wide range of behaviors beyond specific tasks, using an end-to-end learning model that requires no prior preparation such as filtering specific parts of sensory data for the self-prior. Specifically, in the discrete environment model, we learn maximum likelihood estimation via empirical frequency normalization (plug-in MLE), while in the continuous environment model, we train Normalizing Flow based density estimators via maximum likelihood (durkan_neural_2019).
3.1 Model for Discrete Environment
The model for the discrete environment uses categorical distributions as the generative model, handling discrete random variables: